DU TUU UUUUI Slalu! 1. What is the chemical formula for the potential precipitate that might form when a solution of K3PO4 isdo not understand
Answer #1

a) K3PO4 is mixed with Al2(SO4)3 to form the precipitate of AlPO4.

The name of Precipitate is Aluminum phosphate .( AlPO4)

b)

K3PO4 = 15.0 mL of 4.10x10^-3M

Al2(SO4)3 = 9.40mL of 4.35x10^-4 M

Total volume = 15.0+9.40 = 24.4 mL

after mixing

concentration of K3PO4 = 4.10 x10^-3 x 15.0 / 24.4 = 2.52 x10^-3 M

K3PO4 -------------- 3 K+ + PO4-3

concentration of PO4-3 = 2.52x10^-3M = 0.00252M

concentration of Al2(SO4)3 = 4.35x10^-4 x 9.40 / 24.4 = 1.67 x10^-4M

Al2(SO4)3 ------------------ 2 Al+3 + 3 SO4-3

Concentration of Al+3 = 2 x1.67x10^-4 = 3.34x10^-4M

ALPO4(s) -------------------- Al+3(aq) + PO4-3(aq)

Qsp = [Al+3][PO4-3]

Qsp = 3.34x10^-4 x 2.52x10^-3

Qsp = 8.42 x10^-7

c)

Ksp = 9.48x10^-21

Qsp = 8.42x10^-7

Qsp > Ksp

Qsp is greater than Ksp

Precipitate will be form after mixing with K3PO4 and Al2(SO4)3.